OpenStudy (anonymous):

Find the center, vertices, and co-vertices of the ellipse. Then graph the ellipse. ((x^2)/16)) + ((y^2)/25))=1

OpenStudy (mertsj):

Is the major axis horizontal or vertical?

OpenStudy (anonymous):

It doesn't specify.

OpenStudy (mertsj):

You are supposed to know by looking at the equation.

OpenStudy (anonymous):

Oh. I don't

OpenStudy (anonymous):

Oh wait! Is it vertical?

OpenStudy (mertsj):

\[\frac{(x-h)^2)}{b^2}+\frac{(y-k)^2}{a^2}=1\]

OpenStudy (mertsj):

yes.

OpenStudy (mertsj):

That is the equation of an ellipse who major axis is vertical. The center is (h,k) The length of the major axis is 2a The length of the minor axis is 2b The vertices are the endpoints of the major axis The co-vertices are the endpoints of the minor axis

OpenStudy (anonymous):

So, the center would be (0,0) right? And the vertices would be (0,50) and (0,-50)?